# Democratising Agricultural Commodity Price Forecasting: The AGRICAF Approach

### v2

@@ Abstract
Ensuring food security is a critical global challenge, particularly for low-income countries where food prices impact the access to nutritious food. The volatility of global agricultural commodity (AC) prices exacerbates food insecurity, with international trade restrictions and market disruptions further complicating the situation. Despite online platforms for monitoring food prices, there is a significant gap in providing detailed explanations and forecasts accessible to non-specialists. To address this, we propose the Agricultural Commodity Analysis and Forecasts (AGRICAF) methodology, integrating explainable machine learning (XML) and econometric techniques to analyse and forecast global agricultural commodity prices up to one year ahead, dynamically adapting to different forecast horizons. This innovative integration allows us to model complex interactions and dynamics while providing clear, interpretable results. This paper demonstrates how AGRICAF can be used, applying it to three major agricultural commodities—maize, soybean, and wheat—and explaining how different factors impact prices across various months and forecast horizons. By facilitating access to accurate and interpretable medium-term forecasts of AC prices, AGRICAF can contribute to developing a fair and sustainable food system.
@@
Food-security, Agricultural commodity trade, Price forecasting, Food Justice, Explainable Machine learning
@@ Introduction
Changes in the global supply of AC’s are linked to the local Market Year of each area, as the timing of harvesting and market availability in different regions directly impacts the overall global supply chain and commodity prices (FAS-USDA, 2023). Accordingly, the dependent variable in the analysis, $p_{m,y}$, was defined as the proportion (percent) of price change relative to the corresponding month ($m$) of the preceding year ($y-1$), such as $y=1,2,…,Y$. Further details about the calculation process are in Appendix  9.
@@ Results
Fig. 3 presents a comparative analysis of observed and predicted (in sample) global price changes for three distinct agricultural commodities in focus — maize (a), soybean (b), and wheat (c) — relative to twelve forecasting horizons. The temporal dynamics of these commodities are represented over the years 2007 to 2024. For each commodity, the observed relative price change is shown in the black line. The coloured lines show the forecasts obtained by the model with highest accuracy, relative to forecast horizons. Numeric results of the average prediction accuracy are given in Appendix  11 in the form of two error matrices.
@@
Policy Implications
High-income countries often have systems to enhance market transparency and manage price stability through information access, whereas low-income countries frequently experience sharp local price increases during crises, worsening food insecurity due to their limited resources and market information (Headey and Fan, 2008). By offering clear, interpretable forecasts, AGRICAF aims to close gap to support a range of users—policymakers, smallholder farmers, cooperatives, and local community leaders—in better understanding anticipated price movements.
For policymakers, AGRICAF could serve as a resource to help design more responsive food security policies, supporting efforts to stabilise food prices or manage strategic reserves during volatile periods. Smallholder farmers and cooperatives can use AGRICAF’s forecasts to plan planting, storage, and sales, enabling them to plan market risks more effectively. Additionally, by making reliable, data-driven forecasts openly available, AGRICAF supports a more equitable global food system where low-income communities have improved access to essential market information, protecting them from severe price fluctuations that compromise their nutritional and economic well-being (Headey and Hirvonen, 2023; Barrett and Bellemare, 2011).
On a high level, AGRICAF’s accessible price forecasts may contribute to Sustainable Development Goals (SDGs), particularly those focused on food security (SDG 2) and social equity (SDG 10) (UN, 2015). By enabling accessible agricultural commodity price forecasts, AGRICAF can assist policymakers to stabilise price and food supply more effectively, especially in low-income regions prone to food insecurity. With such knowledge, governments can proactively to manage reserves; implement price-stabilisation mechanisms to protect consumers from sudden price shocks; or inform farmers about expected price trends, risks and opportunities for the coming year.
AGRICAF also has implications for smallholder farmers and local community networks, such as NGOs and farmer cooperatives. AGRICAF can empower these agents to make informed decisions on planting, storage, and sales, therefore building greater resilience to price shifts. Smallholders can better navigate the market, optimise and negotiate better income, and reduce risks associated with unfamiliarity with the market conditions (Ferris et al., 2014; K. K. S. Nxumalo and Tekana, 2019). Intermediaries such as government agencies and NGOs can use AGRICAF to extend forecast information to rural areas, thus expanding the reach of these forecasts to farmers who may not have direct access (Adjemian, 2012; Abbott et al., 2016).
In addition to accessibility, AGRICAF emphasises transparency through its XML models, making it more trustworthy for non specialist users (Molnar, 2022) for crisis management and yearly policy planning. With its capacity for monthly updates, policymakers can respond regularly to evolving conditions, such as geopolitical crises or extreme weather events. This adaptability allows AGRICAF to support flexible policy adjustments, offering a tool for managing food supply and price stability in unstable times.
While the long-term policy impacts of AGRICAF remain to be fully evaluated, its accessible approach offers a potentially valuable tool for enhancing market transparency and resilience. By equipping policymakers and smallholder farmers with accessible, data-informed insights, AGRICAF may represent a step forward in addressing the structural challenges that contribute to market instability and food insecurity in low-income settings.
@@ Acknowledgements
Many thanks to Anne Goujon, Sören Lindner, Sebastian Poledna, and Serguei Kaniovski for their valuable feedback, with special appreciation to Anne for her constant advice and support. Christiane Pohn-Hufnagl provided unwavering support beyond research, and I am immensely thankful for her guidance.
I owe much to my PhD advisor, David Makowski, for his continuous support and insightful feedback. Special thanks to Jesus Crespo Cuaresma for his tips and warm welcome into the WU community.
Finally, I appreciate Alan Leavy for his technical help and Clelia Minaudo for her dedicated support as my project adviser.

### v3

@@ Abstract
Ensuring food security is a global challenge, particularly in low-income countries where food prices affect access to nutritious food. The instability of global agricultural commodity (AC) prices exacerbates food insecurity, with international trade restrictions and market disruptions further complicating the situation. Although online platforms exist for monitoring food prices, there is still a need for accessible, detailed forecasts for non-specialists. This paper proposes the Agricultural Commodity Analysis and Forecasts (AGRICAF) methodology, integrating explainable machine learning (XML) and econometric techniques to analyse and forecast global AC’s prices up to one year ahead across different horizons. This innovative integration allows us to model complex interactions while providing clear, interpretable results. We demonstrate the utilization of AGRICAF, applying it to three major commodities and explaining how different factors impact prices across months and forecast horizons. By facilitating access to reliable forecasts of AC prices, AGRICAF can advance a fairer and sustainable food system.
@@
Food-security, Agricultural Commodity Trade, Price Forecasting, Forecasting for Social Good, Explainable Machine Learning
@@ Introduction
Changes in the global supply of AC’s are linked to the local Market Year of each area, as the timing of harvesting and market availability in different regions directly impacts the overall global supply chain and commodity prices (FAS-USDA, 2023). Accordingly, the dependent variable in the analysis, $p_{m,y}$, was defined as the proportion (percent) of price change relative to the corresponding month ($m$) of the preceding year ($y-1$), such as $y=1,2,…,Y$. Further details about the calculation process are in Appendix  8.
@@ Results
Fig. 3 presents a comparative analysis of observed and predicted (in sample) global price changes for three distinct agricultural commodities in focus — maize (a), soybean (b), and wheat (c) — relative to twelve forecasting horizons. The temporal dynamics of these commodities are represented over the years 2007 to 2024. For each commodity, the observed relative price change is shown in the black line. The coloured lines show the forecasts obtained by the model with highest accuracy, relative to forecast horizons. Numeric results of the average prediction accuracy are given in Appendix  10 in the form of two error matrices.
@@ Acknowledgements
Many thanks to Anne Goujon, Sören Lindner, Sebastian Poledna, and Serguei Kaniovski for their valuable feedback, with special appreciation to Anne for her constant advice and support. Christiane Pohn-Hufnagl provided unwavering support beyond research, and I am immensely thankful for her guidance. I owe much to my PhD advisor, David Makowski, for his continuous support and insightful feedback. Special thanks to Jesus Crespo Cuaresma for his tips and warm welcome into the WU community. Finally, I appreciate Alan Leavy for his technical help and Clelia Minaudo for her dedicated support as my project adviser.